jill created a metal containing 28% platnium by combining two other metals one of these other two metals weighed 2 mg and contained 20% platinum if the other weighed 8mg then what percentof it was platinum
1
Expert's answer
2012-09-27T10:38:18-0400
Let second metal contains X parts of platinum. Then

8*X + 2*0.2 = (8 + 2)*0.28,

or

8*X + 0.4 = 2.8

and so, X = 0.4. Therefore, there was 40% of platinum.